# A Fourier-RKHS approach for characterizing equivalent Gaussian distributions for stationary processes on homogeneous spaces

### v1

@@ Abstract
Pairs of equivalent Gaussian distributions for centered stationary processes on homogeneous spaces can be characterized in terms of their spectral measures. The purpose of this note is to consider the latter characterization from the perspective of a reproducing kernel Hilbert space (RKHS) approach.
@@ 1.1 Notational conventions
We refer to $\mathbb{N}$ as the set of strictly positive integers. In the absence of ambiguity, we write $(a_{n})=(a_{n})_{n\in\mathbb{N}}$ for a sequence indexed over $\mathbb{N}$. Given a topological space $(E,\tau)$, the Borel $\sigma$-field over $E$ is denoted by $\mathfrak{B}(E)$. The space of continuous functions $f$ on $E$ is written as $C(E)$. Also, $f\in C_{c}(E)$ if and only if $f\in C(E)$ and $f$ has compact support. For $x,y\in\mathbb{R}^{d}$, $\langle x,y\rangle=x^{\mathrm{t}}y$ identifies the dot product on $\mathbb{R}^{d}$. Given a measure space $(E,\mathcal{A},\mu)$, the space of measurable functions $f\colon E\to\mathbb{C}$ that are square integrable on $E$ w.r.t. $\mu$ is denoted with $L^{2}(E,\mathcal{A},\mu)$. The canonical norm of $\langle f,g\rangle_{\mu}=\int_{E}f(x)\overline{g(x)}\mu(dx)$ on $L^{2}(E,\mathcal{A},\mu)$ is denoted with $\lVert\cdot\rVert_{\mu}$. If clear from the context, we make use of the short notations $L^{2}(\mu)$ or $L^{2}(E)$. The space of absolutely integrable functions (on $E$ w.r.t. $\mu$) is written as $L^{1}(E)$. Given two measures $\mu_{1}$ and $\mu_{1}$ on $\mathcal{A}$ and $L^{2}(\mu_{1})\supset L\subset L^{2}(\mu_{2})$, we use the notation $\lVert\cdot\rVert_{\mu_{1}}\asymp\lVert\cdot\rVert_{\mu_{2}}$ on $L$ to indicate that the norms $\lVert\cdot\rVert_{\mu_{1}}$ and $\lVert\cdot\rVert_{\mu_{2}}$ are equivalent on $L$. That is, there exist constants $\alpha_{1},\alpha_{2}>0$, s.t. for any $\varphi\in L$, $0<\alpha_{1}\lVert\varphi\rVert_{\mu_{2}}\leq\lVert\varphi\rVert_{\mu_{1}}\leq\alpha_{2}\lVert\varphi\rVert_{\mu_{2}}<\infty$. The measures $\mu_{1}$ and $\mu_{2}$ are termed equivalent on $\mathcal{A}$ if they are mutually absolutely continuous on $\mathcal{A}$, i.e., $\mu_{1}(A)=0$ implies $\mu_{2}(A)=0$, $A\in\mathcal{A}$, and vice versa. If $\mu_{1}$ and $\mu_{2}$ are equivalent on $\mathcal{A}$ we write $\mu_{1}\equiv\mu_{2}$ on $\mathcal{A}$. On the other hand, $\mu_{1}$ and $\mu_{2}$ are referred to as orthogonal on $\mathcal{A}$, written as $\mu_{1}\perp\mu_{2}$ on $\mathcal{A}$, if there exists a separating set $A\in\mathcal{A}$ for which $\mu_{1}(A)=0$ and $\mu_{2}(\Omega\setminus A)=0$.
@@ 1.2 Symmetric positive-definite kernels and their RKHS
where $\langle\cdot,\cdot\rangle_{R}$ denotes the inner product on $H_{T}(R)$. This identifies $H_{T}(R)$ as the unique RKHS with reproducing kernel $R$. The existence and uniqueness statement regarding $H_{T}(R)$ is known as the Moore–Aronszajn theorem (cf. [Aronszajn, RKHSBook]). For $n\in\mathbb{N}$ and $T_{n}=\{t_{1},\dotsc,t_{n}\}\subset T$, we write $R(n)$ for the $n\times n$ matrix with entries $R(t_{i},t_{j})$. If $R$ is strictly positive-definite, one can identify $H_{T_{n}}(R(n))=\mathbb{R}^{n}$ with inner product
@@ 1.3 Gaussian processes and their equivalent distributions
It follows that $X$ under $P_{\ell}$ is a Brownian motion with covariance function $R_{\ell}(s,t)=\sigma^{2}\operatorname{min}\{s,t\}$.
A valuable consideration is that $P_{1}$ and $P_{2}$ are either equivalent or orthogonal. This dichotomy has been verified by various authors (cf. [ParzenTimeSeries], p. 478). Feldman [Feldman1958] provides a characterization in terms of the Gaussian spaces $H_{1}(X)$ and $H_{2}(X)$. In particular, he shows that for $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$, it is necessary and sufficient that there exists a linear homeomorphism $T$ from $H_{1}(X)$ onto $H_{2}(X)$ s.t. $T^{*}T-I$ is Hilbert-Schmidt. He refers to $T$ as an equivalence operator from $H_{1}(X)$ onto $H_{2}(X)$. Later, Rozanov [Rozanov] provides an alternative proof using the entropy of the measure $P_{1}$ w.r.t. $P_{2}$. An approach which was pioneered earlier by Hájek [HajekJDiv, HajekOriginal]. Let $T_{n}=\{t_{1},\dotsc,t_{n}\}$ be a finite collection of coordinates from $T$ and denote by $\sigma(Y_{n})$ the $\sigma$-field generated by the Gaussian vector $Y_{n}=(X_{t_{1}},\dotsc,X_{t_{n}})$. Also, let $P_{\ell}^{n}$ be the restriction of $P_{\ell}$ to $\sigma(Y_{n})$ and write $\mathcal{K}$ for the class of all finite subsets of $T$. Then, Hájek [HajekOriginal] shows that
@@
The number (3) is known as the J-divergence (or just divergence) between the finite-dimensional distributions $P_{1}^{n}$ and $P_{2}^{n}$ of $X$ (cf. [Jeffreys] p. 158 and [KullbackLeibler]).
@@
$2J(n)=\operatorname{tr}(A_{n}^{12})+\operatorname{tr}(A_{n}^{21})-2n,\quad A_{n}^{ij}=R_{i}(n)[R_{j}(n)]^{-1},\quad i,j=1,2.$ (6)
If we reconsider Example 1.2 and take $X$ as a Brownian motion with covariance function $R_{\ell}(s,t)=\sigma_{\ell}^{2}\min\{s,t\}$ under $P_{\ell}$ ($\sigma_{1}\neq\sigma_{2}$), it follows from (5) that $P_{1}\perp P_{2}$ on $\sigma_{[0,1]}(X)$. This result can be traced back to the work of Cameron and Martin [CameronMartin] — compare also with Example 1 in Hájek’s paper [HajekOriginal].
@@ 1.4 RKHS characterization of equivalent Gaussian distributions
where for any $k=1,\dotsc,n$, $f_{k}\otimes g_{k}\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ is an elementary tensor. In view of the RKHS choices $H_{T}(R_{1})$ and $H_{T}(R_{2})$, we let $H_{T\times T}(R_{1}\otimes R_{2})$ denote the RKHS on $T\times T$ with kernel $R_{1}\otimes R_{2}\big((s_{1},s_{2}),(t_{1},t_{2})\big)=R_{1}(s_{1},t_{1})R_{2}(s_{2},t_{2})$. Then, it is known that the map $u\mapsto\widetilde{u}$ extends to an isometry between the Hilbert spaces $H_{T}(R_{1})\otimes H_{T}(R_{2})$ and $H_{T\times T}(R_{1}\otimes R_{2})$. This is summarized by writing
$\widetilde{H_{T}(R_{1})\otimes H_{T}(R_{2})}=H_{T\times T}(R_{1}\otimes R_{2}).$ (8)
A proof of the latter equality is given by Aronszajn [Aronszajn] (cf. Section 8). We also refer to Section 5.5 in [RKHSBook]. Early work on the RKHS characterization of equivalent Gaussian distributions was done by Parzen [ParzenRKHS], and Kallianpur and Oodaira [KallianpurOodaira1] (cf. Oodaira [Oodaira]). For instance, for $T$ countably infinite, Parzen [ParzenRKHS] relies on Hájek’s characterization of equivalence (2) and proclaims the necessity of $R_{2}-R_{1}\in H_{T\times T}(R_{1}\otimes R_{2})$ for $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$. Furthermore, an expression of the Radon–Nikodym derivative of $P_{2}$ w.r.t. $P_{1}$ is derived (cf. Capon [Capon]). Later, a general result is obtained by Neveu [Neveu]. For simplicity, we write $H_{T}(R_{\ell})^{2\otimes}=H_{T}(R_{\ell})\otimes H_{T}(R_{\ell})$, $R_{\ell}^{2\otimes}=R_{\ell}\otimes R_{\ell}$ and denote by $H_{T}(R_{\ell})^{2\odot}=H_{T}(R_{\ell})\odot H_{T}(R_{\ell})$ the closed subspace of $H_{T}(R_{\ell})^{2\otimes}$ composed of symmetric tensors. Accordingly, let $R_{\ell}^{2\odot}=R_{\ell}\odot R_{\ell}$ be defined by
@@
In particular, using the isometric correspondence between $H_{T}(R_{\ell})^{2\otimes}$ and $H_{T\times T}(R_{\ell}^{2\otimes})$, it follows that $H_{T}(R_{\ell})^{2\odot}$ and $H_{T\times T}(R_{\ell}^{2\odot})$ are isometric. Let $\langle\cdot,\cdot\rangle_{2\odot}$ denote the inner product on $H_{T}(R_{1})^{2\odot}$ and write $\mathscr{U}\colon H_{T}(R_{1})\to H_{T}(R_{1})^{*}$ for the Hilbert-Schmidt operator $\mathscr{U}(f)(g)=\langle f\odot g,U\rangle_{2\odot}$ associated with an element $U\in H_{T}(R_{\ell})^{2\odot}$. Then, using the inherent structure of the Gaussian space $H_{1}(X)$, Neveu [Neveu] (cf. Proposition 8.6) shows that $P_{1}$ and $P_{2}$ are equivalent on $\sigma_{T}(X)$ if and only if there exists $u\in H_{T\times T}(R_{1}^{2\odot})$ s.t.
@@
and the eigenvalues of the Hilbert-Schmidt operator $\mathscr{U}$, associated with the corresponding element $H_{T}(R_{1})^{2\odot}\ni U$ of $u$, are strictly larger than 1-1. Notice that (9) is equivalent to $R_{2}-R_{1}\in H_{T\times T}(R_{1}^{2\otimes})$. We also point out that for $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$ it is necessary that $H_{T}(R_{1})=H_{T}(R_{2})$. This can be derived from the correspondence $H_{T}(R_{1})=H_{T}(R_{2})$ if and only if $\lVert\cdot\rVert_{P_{1}}\asymp\lVert\cdot\rVert_{P_{2}}$ on $H_{0}(X)$, which is a consequence of Aronszajn’s differences of kernels theorem (cf. [Aronszajn], Corollary IV3 on p. 383). In particular, we can substitute the assumption on the operator $\mathscr{U}$ and obtain the following RKHS characterization of equivalent Gaussian distributions (cf. Chatterji and Mandrekar [Chatterji_Mandrekar]): $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if
@@
That is, $X$ is stationary under $P_{\ell}$. Define $k_{\ell}(t)=R_{\ell}(t,0)$, $t\in\mathbb{R}^{d}$. We observe that $R_{\ell}(s,t)=k_{\ell}(s-t)$. In particular, the assignment $(s,t)\mapsto k_{\ell}(s-t)$ is a symmetric nonnegative-definite kernel. Assume that $k_{\ell}$ is continuous at zero. By Bochner’s theorem (for real coordinate spaces) [Bochner],
@@
In the latter, $h\mapsto\hat{h}$ denotes the Fourier transform of $h$ (cf. Wendland [Wendland_Book], Theorem 10.12). In particular, $H_{\mathbb{R}^{d}}(R_{\ell})\subset L^{2}(\mathbb{R}^{d})$. Upon the isometric correspondence given in (8), it follows that $H_{\mathbb{R}^{d}\times\mathbb{R}^{d}}(R_{\ell}^{2\otimes})\subset L^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d})$. Then, since the mapping
@@
is constant, we conclude that $R_{2}-R_{1}$ does not belong to $H_{\mathbb{R}^{d}\times\mathbb{R}^{d}}(R_{1}^{2\otimes})$, unless $R_{1}$ and $R_{2}$ are equal. This shows that for $T=\mathbb{R}^{d}$, the Gaussian measures $P_{1}$ and $P_{2}$ are orthogonal on $\sigma_{\mathbb{R}^{d}}(X)$ as soon as $R_{1}\neq R_{2}$. A recent treatment — for $T$ not necessarily equal to $\mathbb{R}^{d}$ but sufficiently dense — is given in [hediger]. If $T$ is bounded, the situation is different. As an example, Striebel [Striebel] (cf. [Capon]) gives an expression of the Radon–Nikodym derivative of $P_{2}$ w.r.t. $P_{1}$ for the case where $T=[0,b]$ and
@@
As for a collection of results concerning the characterization of equivalent Gaussian distributions for real stationary processes, an overview is given in the books by Yadrenko [Yadrenko], Ibragimov and Rozanov [GaussianRandomProcesses], and Gikhman and Skorokhod [StochasticProcesses].
@@
Let $T=\mathbb{S}^{d-1}$, $d\geq 3,$ be the unit sphere in $\mathbb{R}^{d}$. Denote with $\Delta\colon\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\to[0,\pi]$ the great-circle (geodesic) distance on $\mathbb{S}^{d-1}$, i.e., $\Delta(s,t)=\operatorname{arccos}(\langle s,t\rangle)$. Assume that
@@
where $\psi_{\ell}\colon[0,\pi]\to\mathbb{R}$ is continuous and s.t. $\psi_{\ell}(0)>0$. Given $k\in\mathbb{N}\cup\{0\}$, write $S_{k}^{j}$, $j=1,\dotsc,h(k)$, for the spherical harmonics of degree $k$ (cf. Chapter XI (Section 11.3) in [Erdelyi] or also Chapter IV (Section 2) in [SteinWeiss]). Then, in the sense of Schoenberg [Schoenberg], the following series representation of $R_{\ell}$ is valid,
@@
In the latter, the coefficients $a_{\ell}(k)$ are strictly positive for infinitely many $k$ (cf. p. 72 in [Yadrenko] or also Theorem 1 in [GneitingMultiquadric]). An explicit description of $H_{\mathbb{S}^{d-1}}(R_{\ell})$ is given in [RKHSSphere]. In particular, any member $f$ of $H_{\mathbb{S}^{d-1}}(R_{\ell})$ is linked with a square summable sequence $(c_{k,j}(f))$, $j=1,\dotsc,h(k)$, s.t.
$f(t)=\sum_{k=0}^{\infty}\sum_{j=1}^{h(k)}S_{k}^{j}(t)c_{k,j}(f)\sqrt{a_{\ell}(k)},\quad\langle f,g\rangle_{R_{\ell}}=\sum_{k=0}^{\infty}\sum_{j=1}^{h(k)}c_{k,j}(f)c_{k,j}(g).$ (11)
Assume w.l.o.g. that $R_{2}-R_{1}\in H_{\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}}(R_{1}^{2\otimes})$ is derived from an elementary tensor, i.e., there exist $f,g\in H_{\mathbb{S}^{d-1}}(R_{1})$ s.t. $(R_{2}-R_{1})(s,t)=f(s)g(t)$. Upon the description given in (11), it follows that
@@
Therefore, $P_{1}\equiv P_{2}$ on $\sigma_{\mathbb{S}^{d-1}}(X)$ if and only if the latter sum is finite. Upon the results from Feldman and Hájek, one can arrive at the same conclusion (cf. Theorem 1 in [GaussianSphereEquivOrtho]).

### v2

# A Fourier-RKHS approach for detecting orthogonal Gaussian distributions for stationary processes on homogeneous spaces

@@ Abstract
Pairs of equivalent Gaussian distributions for centered stationary processes on homogeneous spaces can be characterized in terms of their spectral measures. The purpose of this note is to consider part of the latter characterization from the perspective of a reproducing kernel Hilbert space (RKHS) approach.
@@ 1.1 Notational conventions
We refer to $\mathbb{N}$ as the set of strictly positive integers. In the absence of ambiguity, we write $(a_{n})=(a_{n})_{n\in\mathbb{N}}$ for a sequence indexed over $\mathbb{N}$. Given a topological space $(E,\tau)$, the Borel $\sigma$-field over $E$ is denoted by $\mathfrak{B}(E)$. The space of continuous functions $f$ on $E$ is written as $C(E)$. Also, $f\in C_{c}(E)$ if and only if $f\in C(E)$ and $f$ has compact support. For $x,y\in\mathbb{R}^{d}$, $\langle x,y\rangle=x^{\mathrm{t}}y$ identifies the dot product on $\mathbb{R}^{d}$. Given a measure space $(E,\mathcal{A},\mu)$, the space of measurable functions $f\colon E\to\mathbb{C}$ that are square integrable on $E$ w.r.t. $\mu$ is denoted by $L^{2}(E,\mathcal{A},\mu)$. The canonical norm of $\langle f,g\rangle_{\mu}=\int_{E}f(x)\overline{g(x)}\mu(dx)$ on $L^{2}(E,\mathcal{A},\mu)$ is written as $\lVert\cdot\rVert_{\mu}$. If clear from the context, we make use of the short notations $L^{2}(\mu)$ or $L^{2}(E)$. The space of absolutely integrable functions (on $E$ w.r.t. $\mu$) is identified with $L^{1}(E)$. Given two measures $\mu_{1}$ and $\mu_{2}$ on $\mathcal{A}$ and $L^{2}(\mu_{1})\supset L\subset L^{2}(\mu_{2})$, we use the notation $\lVert\cdot\rVert_{\mu_{1}}\asymp\lVert\cdot\rVert_{\mu_{2}}$ on $L$ to indicate that the norms $\lVert\cdot\rVert_{\mu_{1}}$ and $\lVert\cdot\rVert_{\mu_{2}}$ are equivalent on $L$. That is, there exist constants $\alpha_{1},\alpha_{2}>0$, s.t. for any $\varphi\in L$, $0<\alpha_{1}\lVert\varphi\rVert_{\mu_{2}}\leq\lVert\varphi\rVert_{\mu_{1}}\leq\alpha_{2}\lVert\varphi\rVert_{\mu_{2}}<\infty$. The measures $\mu_{1}$ and $\mu_{2}$ are termed equivalent on $\mathcal{A}$ if they are mutually absolutely continuous on $\mathcal{A}$, i.e., $\mu_{1}(A)=0$ implies $\mu_{2}(A)=0$, $A\in\mathcal{A}$, and vice versa. If $\mu_{1}$ and $\mu_{2}$ are equivalent on $\mathcal{A}$ we write $\mu_{1}\equiv\mu_{2}$ on $\mathcal{A}$. On the other hand, $\mu_{1}$ and $\mu_{2}$ are referred to as orthogonal on $\mathcal{A}$, written as $\mu_{1}\perp\mu_{2}$ on $\mathcal{A}$, if there exists a separating set $A\in\mathcal{A}$ for which $\mu_{1}(A)=0$ and $\mu_{2}(E\setminus A)=0$.
@@ 1.2 Symmetric positive-definite kernels and their RKHS
where $\langle\cdot,\cdot\rangle_{R}$ denotes the inner product on $H_{T}(R)$. This identifies $H_{T}(R)$ as the unique RKHS with reproducing kernel $R$. The existence and uniqueness statement regarding $H_{T}(R)$ is known as the Moore–Aronszajn theorem (cf. [3, 37]). For $n\in\mathbb{N}$ and $T_{n}=\{t_{1},\dotsc,t_{n}\}\subset T$, we write $R(n)$ for the $n\times n$ matrix with entries $R(t_{i},t_{j})$. If $R$ is strictly positive-definite, one can identify $H_{T_{n}}(R(n))=\mathbb{R}^{n}$ with inner product
@@ 1.3 Gaussian processes and their equivalent distributions
It follows that $X$ under $P_{\ell}$ has covariance function $R_{\ell}(s,t)=\sigma^{2}_{\ell}\operatorname{min}\{s,t\}$.
A valuable consideration is that $P_{1}$ and $P_{2}$ are either equivalent or orthogonal. This dichotomy has been verified by various authors (cf. [35], p. 478). Feldman [15] provides a characterization in terms of the Gaussian spaces $H_{1}(X)$ and $H_{2}(X)$. In particular, he shows that for $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$, it is necessary and sufficient that there exists a linear homeomorphism $U$ from $H_{1}(X)$ onto $H_{2}(X)$ s.t. $U^{*}U-I$ is Hilbert-Schmidt. He refers to $U$ as an equivalence operator from $H_{1}(X)$ onto $H_{2}(X)$. Later, Rozanov [38] provides an alternative proof using the entropy of the measure $P_{1}$ w.r.t. $P_{2}$. An approach which was pioneered earlier by Hájek [24, 25]. Let $T_{n}=\{t_{1},\dotsc,t_{n}\}$ be a finite collection of coordinates from $T$ and denote by $\sigma(Y_{n})$ the $\sigma$-field generated by the Gaussian vector $Y_{n}=(X_{t_{1}},\dotsc,X_{t_{n}})$. Also, let $P_{\ell}^{n}$ be the restriction of $P_{\ell}$ to $\sigma(Y_{n})$ and write $\mathcal{K}$ for the class of all finite subsets of $T$. Then, Hájek [25] shows that
@@
The number (3) is known as the J-divergence (or just divergence) between the finite-dimensional distributions $P_{1}^{n}$ and $P_{2}^{n}$ of $X$ (cf. [28] p. 158 and [30]).
@@
$2J(n)=\operatorname{tr}\big[R_{1}(n)R_{2}(n)^{-1}\big]+\operatorname{tr}\big[R_{2}(n)R_{1}(n)^{-1}\big]-2n.$ (6)
If we reconsider Example 1.2 and take $X$ as a Brownian motion with covariance function $R_{\ell}(s,t)=\sigma_{\ell}^{2}\min\{s,t\}$ under $P_{\ell}$ ($\sigma_{1}\neq\sigma_{2}$), it follows from (5) that $P_{1}\perp P_{2}$ on $\sigma_{[0,1]}(X)$. This result can be traced back to the work of Cameron and Martin [8] — compare also with Example 1 in Hájek’s paper [25].
@@ 1.4 RKHS characterization of equivalent Gaussian distributions
where for any $k=1,\dotsc,n$, $f_{k}\otimes g_{k}\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ is an elementary tensor. In view of two RKHS choices $H_{T}(R)$ and $H_{T}(R^{\prime})$, we let $H_{T\times T}(R\otimes R^{\prime})$ denote the RKHS on $T\times T$ with kernel $R\otimes R^{\prime}\big((s_{1},s_{2}),(t_{1},t_{2})\big)=R(s_{1},t_{1})R^{\prime}(s_{2},t_{2})$. Then, it is known that the map $u\mapsto\widetilde{u}$ extends to an isometry between the Hilbert spaces $H_{T}(R)\otimes H_{T}(R^{\prime})$ and $H_{T\times T}(R\otimes R^{\prime})$. This is summarized by writing
$\widetilde{H_{T}(R)\otimes H_{T}(R^{\prime})}=H_{T\times T}(R\otimes R^{\prime}).$ (8)
A proof of the latter equality is given by Aronszajn [3] (cf. Section 8). We also refer to Section 5.5 in [37]. Early work on the RKHS characterization of equivalent Gaussian distributions was done by Parzen [36], and Kallianpur and Oodaira [29] (cf. Oodaira [34]). For instance, for $T$ countably infinite, Parzen [36] relies on Hájek’s characterization of equivalence (2) and proclaims the necessity of $R_{2}-R_{1}\in H_{T\times T}(R_{1}\otimes R_{2})$ for $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$. Furthermore, an expression of the Radon–Nikodym derivative of $P_{2}$ w.r.t. $P_{1}$ is derived (cf. Capon [9]). Later, a general result is obtained by Neveu [33]. For simplicity, we write $H_{T}(R_{\ell})^{2\otimes}=H_{T}(R_{\ell})\otimes H_{T}(R_{\ell})$, $R_{\ell}^{2\otimes}=R_{\ell}\otimes R_{\ell}$ and denote by $H_{T}(R_{\ell})^{2\odot}=H_{T}(R_{\ell})\odot H_{T}(R_{\ell})$ the closed subspace of $H_{T}(R_{\ell})^{2\otimes}$ composed of symmetric tensors. Accordingly, let $R_{\ell}^{2\odot}=R_{\ell}\odot R_{\ell}$ be defined by
@@
In particular, using the isometric correspondence between $H_{T}(R_{\ell})^{2\otimes}$ and $H_{T\times T}(R_{\ell}^{2\otimes})$, it follows that $H_{T}(R_{\ell})^{2\odot}$ and $H_{T\times T}(R_{\ell}^{2\odot})$ are isometric. Let $\langle\cdot,\cdot\rangle_{2\odot}$ denote the inner product on $H_{T}(R_{1})^{2\odot}$ and write $\mathscr{U}\colon H_{T}(R_{1})\to H_{T}(R_{1})^{*}$ for the Hilbert-Schmidt operator $\mathscr{U}(f)(g)=\langle f\odot g,U\rangle_{2\odot}$ associated with an element $U\in H_{T}(R_{1})^{2\odot}$. Then, using the inherent structure of the Gaussian space $H_{1}(X)$, Neveu [33] (cf. Proposition 8.6) shows that $P_{1}$ and $P_{2}$ are equivalent on $\sigma_{T}(X)$ if and only if there exists $u\in H_{T\times T}(R_{1}^{2\odot})$ s.t.
@@
and the eigenvalues of the Hilbert-Schmidt operator $\mathscr{U}$, associated with the corresponding element $H_{T}(R_{1})^{2\odot}\ni U$ of $u$, are strictly larger than 1-1. Notice that (9) is equivalent to $R_{2}-R_{1}\in H_{T\times T}(R_{1}^{2\otimes})$. We also point out that for $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$ it is necessary that $H_{T}(R_{1})=H_{T}(R_{2})$. This can be derived from the correspondence $H_{T}(R_{1})=H_{T}(R_{2})$ if and only if $\lVert\cdot\rVert_{P_{1}}\asymp\lVert\cdot\rVert_{P_{2}}$ on $H_{0}(X)$, which is a consequence of Aronszajn’s differences of kernels theorem (cf. [3], Corollary IV3 on p. 383). In particular, we can substitute the assumption on the operator $\mathscr{U}$ and obtain the following RKHS characterization of equivalent Gaussian distributions (cf. Chatterji and Mandrekar [11]): $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if
@@
That is, $X$ is stationary under $P_{\ell}$. Define $k_{\ell}(t)=R_{\ell}(t,0)$, $t\in\mathbb{R}^{d}$. We observe that $R_{\ell}(s,t)=k_{\ell}(s-t)$. Suppose that $k_{\ell}$ is continuous at zero. Then, by Bochner’s theorem (for real coordinate spaces) [7],
@@
In the latter, $h\mapsto\hat{h}$ denotes the Fourier transform of $h$ (cf. Wendland [46], Theorem 10.12). In particular, $H_{\mathbb{R}^{d}}(R_{\ell})\subset L^{2}(\mathbb{R}^{d})$. Upon the isometric correspondence given in (8), it follows that $H_{\mathbb{R}^{d}\times\mathbb{R}^{d}}(R_{\ell}^{2\otimes})\subset L^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d})$. Then, since the mapping
@@
is constant, we conclude that $R_{2}-R_{1}$ does not belong to $H_{\mathbb{R}^{d}\times\mathbb{R}^{d}}(R_{1}^{2\otimes})$, unless $R_{1}$ and $R_{2}$ are equal. This shows that for $T=\mathbb{R}^{d}$, the Gaussian measures $P_{1}$ and $P_{2}$ are orthogonal on $\sigma_{\mathbb{R}^{d}}(X)$ as soon as $R_{1}\neq R_{2}$. A recent treatment — for $T$ not necessarily equal to $\mathbb{R}^{d}$ but sufficiently dense — is given in [16]. If $T$ is a bounded subset of $\mathbb{R}^{d}$, the situation is different. As an example, Striebel [42] (cf. [9]) gives an expression of the Radon–Nikodym derivative of $P_{2}$ w.r.t. $P_{1}$ for the case where $T=[0,b]$ and
@@
As for a collection of results concerning the characterization of equivalent Gaussian distributions for real stationary processes, an overview is given in the books by Yadrenko [49], Ibragimov and Rozanov [27], and Gikhman and Skorokhod [18].
@@
Let $T=\mathbb{S}^{d-1}$, $d\geq 3,$ be the unit sphere in $\mathbb{R}^{d}$. Denote by $\Delta\colon\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\to[0,\pi]$ the great-circle (geodesic) distance on $\mathbb{S}^{d-1}$, i.e., $\Delta(s,t)=\operatorname{arccos}(\langle s,t\rangle)$. Assume that
@@
where $\psi_{\ell}\colon[0,\pi]\to\mathbb{R}$ is continuous and s.t. $\psi_{\ell}(0)>0$. Given $k\in\mathbb{N}\cup\{0\}$, write $S_{k}^{j}$, $j=1,\dotsc,h(k)$, for the spherical harmonics of degree $k$ (cf. Chapter XI (Section 11.3) in [14] or also Chapter IV (Section 2) in [41]). Then, in the sense of Schoenberg [40], the following series representation of $R_{\ell}$ is valid,
@@
In the latter, the coefficients $a_{\ell}(k)$ are strictly positive for infinitely many $k$ (cf. p. 72 in [49] or also Theorem 1 in [19]). An explicit description of $H_{\mathbb{S}^{d-1}}(R_{\ell})$ is given in [22]. In particular, any member $f$ of $H_{\mathbb{S}^{d-1}}(R_{\ell})$ is linked with a square summable sequence $(c_{k,j}(f))$, $j=1,\dotsc,h(k)$, s.t.
$f(t)=\sum_{k=0}^{\infty}\sum_{j=1}^{h(k)}S_{k}^{j}(t)c_{k,j}(f)\sqrt{a_{\ell}(k)}\quad\text{and}\quad\langle f,g\rangle_{R_{\ell}}=\sum_{k=0}^{\infty}\sum_{j=1}^{h(k)}c_{k,j}(f)c_{k,j}(g).$ (11)
Let $H_{\mathbb{S}^{d-1}}(R_{1})=H_{\mathbb{S}^{d-1}}(R_{2})$ and assume w.l.o.g. that $R_{2}-R_{1}\in H_{\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}}(R_{1}^{2\otimes})$ is derived from an elementary tensor. In particular, there exist $f,g\in H_{\mathbb{S}^{d-1}}(R_{2})$ s.t. $(R_{2}-R_{1})(s,t)=f(s)g(t)$. Upon the description given in (11), it follows that
@@
Therefore, if the latter sum is infinite, the Gaussian measures $P_{1}$ and $P_{2}$ are orthogonal on $\sigma_{\mathbb{S}^{d-1}}(X)$. Upon an explicit evaluation of the divergence (3), one can arrive at the same conclusion (cf. the proof of the necessity part of Theorem 1 in [2]).